OpenStudy (anonymous):

Please double check my answer?? Lidia is packing her suitcase in the dark, so as to not wake her little sister. She blindly chooses 3 of her 5 sweaters. What is the probability that she chooses her 3 favorite sweaters? Option A: 1/10 Option B: 1/15 Option C: 1/60 Option D: 1/75 my answer: option A 1/10

OpenStudy (anonymous):

because I got 3*2*1/5*4*3=6/60=1/10... am i right???

OpenStudy (anonymous):

Hey @ParthKohli again! THanks for coming to help me... again!!! WAs my answer right??

OpenStudy (anonymous):

oops! I meant Hey again @ParthKohli!! hahaa my bad weird phrasing again.. lol :/

Parth (parthkohli):

Favourite or not favourite, so 10 possibilities. S1 S2 S3, let's take these. 5 * 4 * 3 60 since they are not differentiated: 60/3! 60/6 10 Since this is only one combo, then it is 1/10

Parth (parthkohli):

Correct!! ^_^

OpenStudy (anonymous):

YAy!!!! Thank you @ParthKohli you are sooo helpful!! :)

Parth (parthkohli):

Hehe lol